# The Generalization Ridge: Information Flow in Natural Language Generation

### v1

@@ E.2 Incremental Information Gain
In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer. This quantity the information contribution for each transformer blocks.
@@ E.3 Residual Scaling
We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation [12, 13]. Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) [27], Representation-Efficient Fine-Tuning (REFT) [28] and interpretability-driven control [29, 30, 31, 32]. We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers. These values reflect the relative contribution of each layer after optimizing the residual scaling coefficients while keeping all other model parameters frozen. $\beta_{\ell}$ are trained on in-distribution (ID) and out-of-distribution (OOD) data. We observe that in the ID setting, later layers tend to receive higher weights, consistent with memorization behavior. In contrast, OOD training consistently downweights final layers and shifts importance toward the middle of the network, aligning with the generalization ridge identified through InfoRidge.
(a) GPT-2 Small on Synthetic Arithmetic
(b) GPT-2 Small on CLUTRR
(c) Qwen-2.5-0.5B on Synthetic Arithmetic
(d) Qwen-2.5-0.5B on CLUTRR
(e) LLaMA-3.1-8B on Synthetic Arithmetic
(f) LLaMA-3.1-8B on CLUTRR
Figure 10: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. ID training emphasizes later layers, while OOD training shifts weight toward middle layers, aligning with the generalization ridge observed via InfoRidge. Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes 1-sigma error bar.
Appendix F Incremental Information Gain Case Study

### v2

@@ E.2 Incremental Information Gain
In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer. This quantifies the information contribution for each transformer block.
@@ E.3 Residual Scaling
We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024). Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) (Houlsby et al., 2019), Representation-Efficient Fine-Tuning (REFT) (Wu et al., 2024) and interpretability-driven control (Huang et al., 2024; Deng et al., 2025; Meng et al., 2022; Wu et al., 2023). We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers. These values reflect the relative contribution of each layer after optimizing the residual scaling coefficients while keeping all other model parameters frozen. $\beta_{\ell}$ are trained on in-distribution (ID) and out-of-distribution (OOD) data. We observe that in the ID setting, later layers tend to receive higher weights, consistent with memorization behavior. In contrast, OOD training consistently has lower final layers’ weight compared to the ID settings and shifts importance toward the intermediate layers. When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 7. The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”. Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models. In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
(a) GPT-2 Small on CLUTRR
(b) GPT-2 Small on Synthetic Arithmetic
(c) Qwen-2.5-0.5B on CLUTRR
(d) Qwen-2.5-0.5B on Synthetic Arithmetic
(e) LLaMA-3.1-8B on CLUTRR
(f) LLaMA-3.1-8B on Synthetic Arithmetic
Figure 9: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. ID training emphasizes later layers, while OOD training has lower final layer weight. Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes 1-sigma error bar.
Appendix F Attention Dynamics Across Layers Case Study
Appendix G Incremental Information Gain Case Study
@@ G.3 Implication
Appendix H Activation Patching Experiment
To assess which residual blocks causally support OOD prediction, we conduct a targeted activation-patching experiment. Let the model contain $L$ residual blocks with hidden states $Z_{\ell}\in\mathbb{R}^{d}$, and let $t^{*}$ denote the final context token (the last token before next-token prediction). For every OOD sequence $x_{\mathrm{ood}}$, we pair it with an ID sequence $x_{\mathrm{id}}$ sharing the same token structure. For each depth $\ell\in\{1,\dots,L\}$, we compute two forward passes: a standard OOD pass and a patched pass in which we replace only the hidden state of token $t^{*}$ at depth $\ell$ according to $Z_{\ell}(x_{\mathrm{ood}},t^{*})\leftarrow Z_{\ell}(x_{\mathrm{id}},t^{*}).$ All other activations remain unchanged. Let $logit_{\mathrm{base}}$ denote the logit assigned to the correct next token under the baseline OOD pass, and let $logit_{\mathrm{patch}}^{(\ell)}$ denote the corresponding logit under layer-$\ell$ patching. We measure causal impact using the logit drop $\Delta_{\ell}=logit_{\mathrm{base}}-logit_{\mathrm{patch}}^{(\ell)},$ where a positive value indicates that depth $\ell$ provides information necessary for correct OOD prediction.
Table 5 reports the averaged logit drop for all $\ell=1,\dots,12$ (GPT-2 Small on Synthetic Dataset). The effects are negligible for layers 1–8, begin rising around layer 9, increase sharply at layer 10, peak at layer 11, and decline again at layer 12. This rise–peak–fall pattern closely matches the ridge structure observed in Figure 2, where $I(Z_{\ell};Y)$ increases through mid-to-late depth, reaches a maximum in the ridge region, and then decreases. The activation-patching experiment therefore provides supporting causal evidence that the ridge layers identified by IndoRidge are necessary for successful OOD generalization.
Table 5: Average logit drop $\Delta_{\ell}$ (baseline logit - patched logit). Larger values indicate greater causal contribution to OOD prediction.
Layer $\ell$ 1 2 3 4 5 6 7 8 9 10 11 12
$\Delta_{\ell}$ -0.1037 -0.1039 -0.0327 -0.5405 -5.1663 -8.3992 -5.8914 -15.0996 -3.0998 9.3148 17.2022 15.5943
Appendix I Extending Predictive Information to Multi-Token Outputs
To complement our single-token analysis, we further evaluate information flow under multi-token outputs using the CNN/Daily Mail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015).
Autoregressive next-token information.
We follow the model’s natural decoding process. We analyze $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ since there is no deterministic gold answer for each intermediate step in NLG, where for each generation step $t$, $Z_{\ell}^{(t)}$ is the hidden state of current token at layer $\ell$ and $\hat{Y}^{(t+1)}$ is the embedding of model’s predicted token at step $t+1$.
Across steps, the pattern is consistent (Figure 11): information rises from early layers, peaks in the middle layers, and then declines toward top layers. This gives a clear information ridge, indicating that intermediate layers encode the most stable and task-relevant core information that supports the model’s decision-making (generalization). In particular, high $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ at these layers means the model’s predictive direction is most strongly determined by that layer: layers whose representations consistently constrain the next-token prediction across many contexts must encode the shared, invariant structure that the model relies on to generalize. In contrast, upper layers exhibit behavior that might be related to pretrained surface patterns from corpus such as token co-occurrence, causing the representation to drift away from this decision-relevant structure (memorization).
The first step is a notable outlier: it is the only step that conditions solely on the input prompt, and shows the sharpest and most pronounced ridge. Generating the first token requires a global transition from the prompt into the answer trajectory, which heavily engages the abstract, decision-relevant subspace in the middle layers. Crucially, this is also the point where information compression is most extreme, the model must condense all prompt-level semantics into a single initial token decision, therefore forms a substantial ridge. Once generation begins, later generation steps move into a more regular autoregressive regime, where each hidden state already contains a compressed summary of all previously generated content, and the continuation depends more on local context than on global restructuring. Consequently, the depth profiles become flatter and more stable across steps, although the mid-layer ridge remains visible.
(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
Figure 11: $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
Pooled-answer information.
We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y^{\text{pool}})$, where $Y^{\text{pool}}$ is a fixed target vector obtained by average-pooling the embedding-layer representations of all ground-truth answer tokens. Across models and steps (Figure 12), a consistent pattern emerges: intermediate layers are where the hidden state is most strongly aligned with the gold-answer.
(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
Figure 12: $I(Z_{\ell}^{(t)};Y^{\text{pool}})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
Taken together, the two constructions reveal a coherent picture of depth specialization in multi-token generation. Intermediate layers are simultaneously the most informative about the model’s next-token decisions and the most aligned with the global answer. Thus, the two observed ridges are complementary projections of a single underlying phenomenon: intermediate layers concentrated most meaningful and decision-relevant information. In contrast, upper layers increasingly reflect pretrained corpus-specific surface patterns, causing the representation to drift away from this decision-relevant structure.
(a) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
(b) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
Figure 13: Mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI.
We also plot the mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI (Figure 13). These tokens are more predictable because they follow strict grammatical and structural patterns, whereas content words depend more on semantic context and therefore exhibit greater variability.
Appendix J Non-finetuned scenario
Beyond fine-tuning scenarios, we also computed predictive-information $I(Z_{\ell};Y)$ curves in a non-training zero-shot regime and observed the same mid-layer information peak pattern, see Figure 14.
Figure 14: GSM8K on Qwen2.5-0.5B shows an intermediate-layer MI rise and late-layer decline.
Appendix K Kernel Ablation
The matrix-based MI estimator requires a positive-definite Gram matrix, for which the Gaussian kernel is the standard choice. To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels. While these kernels introduce quantitative shifts in the magnitude, the trend—including the rise–peak–decline structure and the location of the ridge—remains unchanged (Figure 15).
Polynomial Kernel.
Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as
$(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ (4)
where $p$ is the polynomial degree and $c_{0}$ is a constant bias term.
Laplacian Kernel.
Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as
$(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ (5)
where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
Figure 15: Different kernel shows similar pattern.
Appendix L Use of Large Language Models.
Large language models (LLMs) were used solely as assistive tools for proofreading and improving clarity of writing.